Charlies Furniture Store has been in business for several years. The firms owners have described the store as a high-price
Answer #1

A. Current sales = turnover*total assets

= 0.2*1,200,000 =$240,000

ROI = Margin*turnover

= 34%*0.2

= 6.8%

B.turnover required = 6.8%/20% = 0.34

Sales required = 0.34*1,200,000 =$408,000

C.Actual increase in sales required =408,000-240,000

=$168,000

I.e. 168,000/240,000

= 70%

D.when the prices are not reduced, margin will only be reduced by the amount spent on advertising since margin = net operating income/sales and advertising cost will lead to a reduction in net operating income

Turnover might increase based on campaign as it will lead to higher sales

ROI will not be affected much, it might reduce a little because of cost of campaign

E. Alternative strategy is to reduce operating assets